OpenStudy (lilai3):

Solve. h^2 +15 = -16h WILL REWARD MEDAL/FAN!!

OpenStudy (tkhunny):

Have you considered factoring?

OpenStudy (lilai3):

h^2 + 15 +16h =0

OpenStudy (tkhunny):

Okay, usually h^2 + 16h + 15 = 0, but let's go with it. Now (h + ____)(h + ____) = 0 Fill in the blanks. Factors of +15 that add to +16. What could it be...

OpenStudy (lilai3):

1 and 15

OpenStudy (tkhunny):

Perfect. Put them in the spaces and tell me what's next.

OpenStudy (lilai3):

(h+15) (h+1) =0 h could equal to -15 or -1.

OpenStudy (tkhunny):

Rock on. Do you believe or are you still doubting?

OpenStudy (lilai3):

Uh, I understand Thank you, @tkhunny